Determine the magnitude of the line current that flows to a load that consumes 3MW at a power factor of 0.8 overdue and that operates at 13.2 kV voltage
Answer #1

Here the type of the system is not given either single phase or 3-phase. Usually 13.2kv exists in 3-phase supply.

1) Assuming it as 3-phase supply :-

Given :-

Power = 3MW

Power factor = 0.8 ( since, magnitude of current is asked, there is no need of lag/ lead. Answer is same for both )

Line Voltage = 13.2kv

We know that P= √3 * V​​​​​​​​​​​​​​L*I​​​L*PF

From that , I​​​​​​L = P/( √3 * V​​​​​​L *PF)

I​​​​​​L = 3M/(√3 * 13.2k * 0.8 )

I​​​​​​L = 164.02 A.

B) Assuming it as a single phase supply:-

Power = V*I*PF

Current I = P/( V*PF)

I = 3M/( 13.2k * 0.8 )

I = 284.09 A

Usually 13.2kv supply is a exists in India as three-phase supply.